# Enhancing Speech Large Language Models with Prompt-Aware Mixture of Audio Encoders

### v1

@@ 3.2 Model Architecture and Training
We train our model based on Huggingface Transformers Library [^1] https://github.com/huggingface/transformers. Our model consists of three audio encoders, a pre-fusion adapter for each encoder, a PaM fusion module, and an LLM. In our main experiments, the encoders are Whisper-Small Radford et al. (2022), WavLM-Base-Plus Chen et al. (2022), and Wav2Vec2-Base-960h Baevski et al. (2020), each with approximately 100 million parameters. We downsample the features from the Whisper encoder by a factor of two, resulting in a frame length of 40ms, consistent with the frame length of Wav2Vec2 and WavLM. The pre-fusion adapter is an FFN that transforms the encoder’s hidden dimension $D_{\text{E}}$ to the LLM’s hidden dimension $D_{\text{LLM}}$. Each expert in the PaM fusion module includes a fusion weight matrix ($\mathbb{R}^{3L\times 3}$) and a linear layer ($\mathbb{R}^{6D_{\text{LLM}}\times D_{\text{LLM}}}$) to fuse features from all encoders. Here, $L$ represents the number of layers in the encoder, which is 12 for all encoders in our experiments. For the fused features, we set $k=3$, corresponding to the number of last hidden states. We utilize four routed experts, each corresponding to a specific task category: ASR-clean, ASR-noisy, SNV, and AC. For each category, we generate 50 prompts using ChatGPT OpenAI (2023). For the LLM model, we select the Qwen2.5-3B Team (2024). In section 4, we also experiment with other encoders, including Hubert-Base-LS960, Whisper-Large-v3, and WavLM-Large.
We train our model for five epochs with a learning rate of 5e-5, 2000 warmup steps, and bf16 precision. We freeze all encoders and the LLM, training only the added parameters of the adapters and the fusion modules. For the LLM, we apply LoRA Hu et al. (2022) with a rank of 32 and an alpha of 64, adding LoRA only on the q_proj and k_proj. Each task has the same probability during training. During the inference stage, we select the last checkpoint on the validation set and perform greedy search.
@@ 4.1 Main Results
As demonstrated in Table 2, we compare the proposed PaM method with single and multi-encoder baselines. Each encoder exhibits distinct advantages. When utilizing a single encoder, the Speech LLM with the Whisper encoder performs best on the AMI dataset and AC tasks. The primary reason is that the Whisper model is trained on a vast amount of speech data, exposing it to diverse acoustic conditions. Consequently, it excels in challenging ASR and AC tasks in real-world environments and noisy conditions. The WavLM encoder, trained on multi-speaker speech signals, provides the best features for the SNV task. The Wav2Vec2 encoder performs best on the LibriSpeech dataset mainly because it was pretrained on this dataset. However, since the LibriSpeech dataset consists of clean audiobooks, the Speech LLM with the Wav2Vec2 encoder shows poor performance on the AMI and AudioCap datasets.
We reimplemented the feature fusion methods of WavLLM Hu et al. (2024) and SALMONN Tang et al. (2024), training the Speech LLM with the three audio encoders in our setups. Both methods use concatenation but are followed by different projection layers: WavLLM with a linear layer and SALMONN with a Q-former layer. Additionally, we implemented a simple averaging method that directly computes the average of the features from the three encoders. Compared to the best performance of single encoder baselines, all three fusion methods achieve better METEOR scores on AC tasks. However, performance may degrade on other tasks. For example, we observed performance degradation for all three methods on the LibriSpeech test-clean subset. This is expected since the same features are used for all tasks. Features containing more useful acoustic information for AC tasks may lack useful semantic information for ASR tasks.
@@ 4.2 Feature Importance
In Figure 3, we visualize the fusion weights for each expert, excluding the shared expert, which can be interpreted as the fusion weight for each task. To enhance clarity, we summed the weights for every four layers, resulting in the total weight for shallow, middle, and deep layers. Generally, different tasks require different features, so each expert has distinct fusion weights. Specifically, when the expert for ASR-clean is activated, it mainly focuses on features from WavLM and Wav2Vec2, especially the deep layers. When the expert for ASR-noise is activated, it primarily focuses on features from the Whisper encoder and WavLM. For SNV and AC tasks, all three encoders have similar fusion weights. For the SNV task, features from the middle layers are more important, while for the AC task, features from the shallow layers are more important.
@@ 4.3 Combinations of Encoders
In Table 3, we extend our investigation to encompass more combinations of encoders, including two and three encoders, and incorporate the HuBERT encoder. To highlight the strengths of each encoder, we calculate the performance by keeping one encoder fixed and varying the other encoders, then computing the average. The average (AVG) rank reflects the overall multitask performance. It is evident that using three encoders significantly outperforms using two encoders in all combinations, thereby demonstrating the effectiveness of employing more encoders for Speech LLMs.
We can observe that different encoders offer varying benefits, which proves that the tasks-specific encoders could significantly improve the performance on corresponding tasks. For example, when Whisper is used, regardless of whether two or three encoders are employed, the Speech LLM achieves the lowest WER on AMI and the highest METEOR scores on AudioCaps. On the other hand, Wav2Vec2 provides an advantage for recognizing speech signals in LibriSpeech. This indicates that when selecting encoders for Speech LLM, it is essential to consider the domain, downstream tasks, and the capabilities of each encoder. It is suggested to use a robust general domain model like Whisper in combination with domain-specific encoders such as Wav2Vec2.
Models LibriSpeech AMI SNV AudioCaps AudioCaps QA
@@ 4.4 Larger Encoders and LLMs
In our investigation of other LLMs, including Qwen2.5-7B, LLaMA3.2-3B, and LLaMA3.1-8B, we observed some improvements in certain tasks. However, the overall performance was not superior to that of Qwen2.5-3B. The potential reason for this is that we used short audios, and both the prompts and answers were brief, thereby not fully utilizing the strong semantic understanding capabilities of the larger LLMs. Consequently, we opted to use Qwen2.5-3B in this paper. It is important to emphasize that for Speech LLMs, the extracted features may be more critical than the LLM itself for many downstream tasks.
@@ 4.5 Parameters of the Adapter
In PaM, we employ multiple experts, merge and concatenate various features. Consequently, the number of parameters is slightly higher than that of the baseline Concatenation and Average methods. To ensure a fair comparison, we reduce the dimensionality within PaM, resulting in a parameter count of only 29M, which is similar to the baselines. As observed in Table 4, PaM outperforms or performs comparably to the baselines across all tasks.
@@ 5 Discussions
More and Unseen Tasks: Although our experiments involve three tasks and five datasets, they are representative as they encompass both semantic and acoustic-related tasks. PaM can be extended to other tasks, such as speech translation, which requires features similar to ASR and thus is not included in our paper. Additionally, PaM is applicable to unseen tasks and datasets. To achieve this, tasks must be categorized based on similar feature requirements, and routing during inference should be based on these categories.
Alignment: Although our paper primarily emphasizes the role of PaM in feature fusion, it also serves an alignment function by mapping the different fused features into the LLM’s embedding space. Since we use different features for different prompts/tasks, the projections needed for feature aligment also vary. This is why each expert not only has fusion parameters but also its own FFN for projection.
@@ 6 Related Works
Audio encoders transform raw waveforms or Fbank features into a high-level feature space that is informative for downstream tasks. These encoders can be classified into two categories: supervised models and self-supervised models. Supervised models typically employ ASR tasks to train an end-to-end model that includes an audio encoder and a text decoder, for instance, an attention-based decoder or a Connectionist Temporal Classification (CTC) decoder. By omitting the decoder, the encoder can serve as a feature extractor Radford et al. (2022); Baevski et al. (2020). Self-supervised models can be trained on unlabeled speech signals. For instance, Wav2Vec2 Baevski et al. (2020) and HuBERT Hsu et al. (2021) were trained to predict the pseudo-discrete target at masked time steps. WavLM Chen et al. (2022) is a variant of HuBERT, designed to facilitate speaker identity extraction by using multi-speaker signals. Different model architectures, training methods, and data can result in encoders with distinct properties and advantages, making the mixture of audio encoders effective for Speech LLMs.
@@
The native LLM handles only text modality. For the LLM to directly process audio modality, it must comprehend audio features, which contain significantly more information than transcriptions and are substantially longer than text embeddings. Since text is represented as discrete tokens, it is natural to extract discrete tokens from continuous speech signals and then expand the vocabulary of text LLMs to understand these speech tokens Rubenstein et al. (2023); Veluri et al. (2024); Ma et al. (2024). An alternative is to use an adapter layer, to directly convert the continuous speech features to the continuous embedding space of the LLM. For example, QwenAudio Chu et al. (2024) employs average pooling to downsample speech features, followed by two linear layers for projection. SALMONN Tang et al. (2024) utilizes the Q-former Yu et al. (2024), a cross-attention-based adapter, to achieve a higher compression ratio. Compared to previous works, our adapter handles more encoders and generates different features based on the prompt, rather than a single feature for all prompts.
@@
With the introduction of the Transformer architecture Vaswani et al. (2017), increasing model size has been shown to improve performance significantly He et al. (2016); Kaplan et al. (2020). This has led to growing interest in the MoE approach, which replaces the FFN sub-layer in Transformer models with multiple experts, thereby enabling substantial parameter growth while maintaining constant inference costs Shazeer et al. (2017). These MoE methods typically employ massive experts and extremely sparse activation routing, without explicitly considering the specialization of individual experts Fedus et al. (2022); Lepikhin et al. (2020). However, the vast scale of these models presents significant deployment challenges. In contrast, the earliest MoE research introduced a data-dependent, trainable combining method Jacobs et al. (1991); Masoudnia and Ebrahimpour (2014), which aims to decompose complex tasks into simpler sub-tasks, each managed by a dedicated expert. Such works have inspired recent advances in developing modular models with emphasis on handling each task by a task-specific expert called expert specialization Ma et al. (2018); Gupta et al. (2022), providing solutions for deploying large-scale MoE models Lu et al. (2024) and enabling individual experts to learn and decompose diverse knowledge Dai et al. (2024). Inspired by these insights, we proposed a specialized MoE fusion method integrating multiple audio features to enhance Speech LLMs.
@@ 7 Conclusion
In conclusion, we propose PaM, a MoE-based feature fusion method designed to provide Speech LLM with diverse features from multiple encoders based on users’ input prompts. Experimental results indicate that PaM surpasses both single-encoder and multi-encoder baselines across a variety of tasks and datasets. We provide a detailed analysis of the feature importance of different encoders, demonstrating that PaM effectively leverages different encoders and levels of features for distinct tasks. Additionally, we present comprehensive experimental results for the selection and combination of encoders. For future work, we intend to expand the training data and incorporate additional tasks.
@@ Appendix A Appendix
A.1 Details of Models and Datasets
In this paper, we leverage multiple audio encoders and LLM to construct the end-to-end speech LLM. Our training dataset is sourced from commonly used open-source datasets, totalling approximately 450 hours of audio data, corresponding to 313,208 samples, as outlined in Table 5.
Table 5: The whole training datasets.
@@ A.1 Details of Models and Datasets
In our paper, we adopt multiple pre-trained audio encoders and LLMs, we list the architecture setting for all model we used in our experiments in Table 6. Notably, for the Whisper model, we only used its encoder part as an audio feature extractor.
@@
Table 6: The settings of pre-trained model we used in our experiments. For the audio encoder models, we utilize only the encoder component and freeze all parameters. For the LLMs, we freeze the base model parameters and apply LoRA adapters to fine-tune the model.

### v2

@@ 3.2 Model Architecture and Training
We train our model based on Huggingface Transformers Library [^4] https://github.com/huggingface/transformers. Our model consists of three audio encoders, a pre-fusion adapter for each encoder, a PaM fusion module, and an LLM. In our main experiments, the encoders are Whisper-Small encoder, WavLM-Base-Plus, and Wav2Vec2-Base-960h [^5] The links to the pretrained models and datasets used can be found in Appendix A.2. The implementation details of baseline methods are available in Appendix A.3., each with approximately 100 million parameters. We downsample the features from the Whisper encoder by a factor of two, resulting in a frame length of 40ms, consistent with the frame length of Wav2Vec2 and WavLM. The pre-fusion adapter is an FFN that transforms the encoder’s hidden dimension $D_{\text{E}}$ to the LLM’s hidden dimension $D_{\text{LLM}}$. Each expert in the PaM fusion module includes a fusion weight matrix ($\mathbb{R}^{3L\times 3}$) and a linear layer ($\mathbb{R}^{6D_{\text{LLM}}\times D_{\text{LLM}}}$) to fuse features from all encoders. Here, $L$ represents the number of layers in the encoder, which is 12 for all encoders in our experiments. For the fused features, we set $K=3$, corresponding to the number of last hidden states. We utilize four routed experts, each corresponding to a specific task category: ASR-clean, ASR-noisy, SNV, and AC. For each category, we generate 50 prompts using ChatGPT OpenAI (2023) [^6] We provide few examples of the prompts we used in Appendix A.11. For the LLM model, we select the Qwen2.5-3B Team (2024). In section 4, we also experiment with other encoders, including Hubert-Base-LS960, Whisper-Large-v3, and WavLM-Large. [^7] Additionally, we add the BEATs model, which performs well on the AC task, to further enhance PaM in Appendix A.5.
We list the training and inference parameters in Appendix A.6.
@@ 4.1 Main Results
As demonstrated in Table 1, we compare the proposed PaM method with single and multi-encoder baselines. Each encoder exhibits distinct advantages. When utilizing a single encoder, the Speech LLM with the Whisper encoder performs best on the AMI dataset and AC tasks. The primary reason is that the Whisper model is trained on vast speech data, exposing it to diverse acoustic conditions. Consequently, it excels in challenging ASR and AC tasks in real-world environments and noisy conditions. The WavLM encoder, trained on multi-speaker speech signals, provides the best features for the SNV task. The Wav2Vec2 encoder performs best on the LibriSpeech dataset mainly because it was pretrained on this dataset. However, since the LibriSpeech dataset consists of clean audiobooks, the Speech LLM with the Wav2Vec2 encoder shows poor performance on the AMI and AudioCap datasets.
We reimplemented the feature fusion methods of WavLLM and SALMONN, training the Speech LLM with the three audio encoders in our setups. Both methods use concatenation but are followed by different projection layers: WavLLM with a linear layer and SALMONN with a Q-former layer. Additionally, we implemented a simple averaging method that directly computes the average of the features from the three encoders. Compared to the best performance of single encoder baselines, all three fusion methods achieve better METEOR scores on AC tasks. However, performance may degrade on other tasks. For example, we observed performance degradation for all three methods on the LibriSpeech test-clean subset. This is expected since the same features are used for all tasks. Features containing more useful acoustic information for AC tasks may lack useful semantic information for ASR tasks.
@@ 4.2 Feature Importance
Figure 3: The weights for each encoder and its layers.
In Figure 3, we visualize the fusion weights for each expert, excluding the shared expert, which can be interpreted as the fusion weight for each task. We summed the weights for every four layers to enhance clarity, resulting in the total weight for shallow, middle, and deep layers. Generally, different tasks require different features, so each expert has distinct fusion weights. Specifically, when the expert for ASR-clean is activated, it mainly focuses on features from WavLM and Wav2Vec2, especially the deep layers. When the expert for ASR-noise is activated, it primarily focuses on features from the Whisper encoder and WavLM. For SNV and AC tasks, all three encoders have similar fusion weights. For the SNV task, features from the middle layers are more important, while for the AC task, shallow layers contribute more.
@@ 4.3 Combinations of Encoders
In Table 2, we extend our investigation to encompass more combinations of encoders, including two and three encoders, and incorporate the HuBERT encoder. To highlight the strengths of each encoder, we calculate the performance by keeping one encoder fixed and varying the other encoders, then computing the average. The average (AVG) rank reflects the overall performance across multitasks. It is evident that using three encoders significantly outperforms using two encoders in all combinations, thereby demonstrating the effectiveness of employing more encoders for Speech LLMs.
We can observe that different encoders offer varying benefits, which proves that the task-specific encoders can significantly improve the performance on the corresponding tasks. For example, when Whisper is used, regardless of how many encoders are employed, the Speech LLM achieves the lowest WER on AMI and the highest METEOR scores on AudioCaps. On the other hand, Wav2Vec2 provides an advantage for recognizing speech signals in LibriSpeech. This indicates that when selecting encoders for Speech LLM, it is essential to consider the domain, downstream tasks, and the capabilities of each encoder. It is suggested to use a robust general domain model like Whisper in combination with domain-specific encoders such as Wav2Vec2.
Models
LibriSpeech AMI SNV AudioCaps AudioCaps QA
@@ 4.4 Larger Encoders and LLMs
In our investigation of other LLMs, including Qwen2.5-7B, LLaMA3.2-3B, and LLaMA3.1-8B, we observed some improvements in certain tasks. However, the overall performance was not superior to that of Qwen2.5-3B. The potential reason for this is that we used short audios, and both the prompts and answers were brief, thereby not fully utilizing the strong semantic understanding capabilities of the larger LLMs. Consequently, we opted to use Qwen2.5-3B in this paper. It is important to emphasize that for Speech LLMs, the extracted features may be more critical than the LLM itself for many downstream tasks. In addition, we also compare the concatenation fusion strategy (WavLLM) and PaM using larger LLMs based on the LLaMA3.1-8B and Qwen2.5-7B. We find that the performance of the concatenation fusion strategy is inconsistent across different base models of similar size, whereas PaM maintains stability, as detailed in Appendix A.7.
@@ 4.5 Parameters of the Adapter
In PaM, we employ multiple experts, merge and concatenate various features. Consequently, the number of parameters is slightly higher than that of the baseline Concatenation and Average methods. To ensure a fair comparison, we reduce the dimensionality within PaM, resulting in only 29M total parameters, similar to the baselines. PaM outperforms the baseline across almost all tasks, with similar overall parameters. Notably, during inference, PaM activates only 26M parameters, in contrast to the 37M parameters activated by the concatenation method, demonstrating the efficiency of PaM. In this configuration, each expert contains only 0.9M parameters, which is smaller than other components of the model, such as the LLM and encoders. Consequently, PaM can be further enhanced by increasing the number of experts with minimal impact on computational cost.
@@ 5 Discussions
Ablation study and routing method: To further validate the effectiveness of PaM, we conducted an ablation study, like PaM without the shared expert, as detailed in Appendix A.8. We also compare PaM against a learnable routing approach without task information, commonly employed in MoE models. The results indicate that PaM is more efficient and better suited to handling multiple downstream tasks than conventional routing and fusion strategies. Moreover, Table 3 highlights the strengths of each encoder. These findings encourage researchers to select encoders tailored to their downstream tasks for the Audio-LLM. While we did not incorporate this prior knowledge into the routing mechanism, leveraging it presents a promising direction for future work.
Other further work: 1.More and Unseen Tasks. Although our experiments involve three tasks and five datasets, they are representative as they encompass both semantic and acoustic-related tasks. We believe PaM can be extended to other tasks, which we will validate in future work. 2.Leverage Multimodal LLMs. Features across various encoders hinder initialization from pretrained multimodal LLMs in our work (Appendix A.9). We will explore strategies to more effectively leverage pretrained multimodal LLMs, such as Lai et al. (2024). 3.Efficient. Improving the efficiency of LLMs has attracted much attention. Enhancing the computational efficiency of speech LLMs presents another promising avenue for exploration (Appendix A.9).
@@ 6 Related Works
Audio encoders can be classified into supervised and self-supervised models. Supervised models typically employ ASR tasks to train an end-to-end model with an audio encoder and a text decoder. By omitting the decoder, the encoder can serve as a feature extractor Radford et al. (2022); Baevski et al. (2020). Self-supervised models can be trained on unlabeled speech signals. For instance, Wav2Vec2 Baevski et al. (2020) and HuBERT Hsu et al. (2021) were trained to predict the pseudo-discrete target at masked time steps. WavLM Chen et al. (2022) is a variant of HuBERT, designed to facilitate speaker identity extraction by using multi-speaker signals. Different model architectures, training methods, and data can result in encoders with distinct properties and advantages, making the mixture of audio encoders effective for Speech LLMs.
@@
To construct end-to-end speech LLMs, a natural approach is to extract discrete tokens from continuous speech signals and then expand the vocabulary of text LLMs to understand these speech tokens Rubenstein et al. (2023b); Veluri et al. (2024); Ma et al. (2024a). An alternative is to use an adapter layer to directly convert the continuous speech features into the continuous embedding space of the LLM. For example, QwenAudio Chu et al. (2024) employs average pooling to downsample speech features, followed by two linear layers for projection. SALMONN Tang et al. (2024) utilizes the Q-former Yu et al. (2024), a cross-attention-based adapter, to achieve a higher compression ratio. In parallel, to achieve high compression, Soundwave replaces cross-attention with a lightweight self-attention module that treats the inputs as query Zhang et al. (2025). Compared to previous works, our adapter handles more encoders and generates different features based on the prompt, rather than a single feature for all prompts.
@@
MoE has attracted growing interest, which replaces the FFN sub-layer in Transformer models with multiple experts Shazeer et al. (2017). These MoE methods typically employ massive experts and extremely sparse activation routing, increasing model size while maintaining constant inference costs, without explicitly considering the specialization of individual experts Fedus et al. (2022); Lepikhin et al. (2020). However, the vast scale of these models presents significant challenges for deployment. In contrast, the earliest MoE research introduced a data-dependent, trainable combining method Jacobs et al. (1991); Masoudnia and Ebrahimpour (2014), which aims to decompose complex tasks into simpler sub-tasks, each managed by a dedicated expert. Such works have inspired recent advances in developing modular models called expert specialization Ma et al. (2018); Gupta et al. (2022), providing solutions for deploying large-scale MoE models Lu et al. (2024) and enabling individual experts to learn and decompose diverse knowledge Dai et al. (2024). Inspired by these insights, we proposed a specialized MoE fusion method integrating multiple audio features to enhance Speech LLMs.
@@ 7 Conclusion
In conclusion, we propose PaM, a feature fusion method designed to provide Speech LLM with diverse features from multiple encoders based on users’ input prompts. Experimental results indicate that PaM surpasses both single-encoder and multi-encoder baselines across a variety of tasks and datasets. We provide a detailed analysis of the feature importance of different encoders, demonstrating that PaM effectively leverages different encoders and levels of features for distinct tasks. Additionally, we present comprehensive experimental results for the selection and combination of encoders. For future work, we intend to expand the training data and incorporate additional tasks.
@@ 8 Limitations
Acknowledgements
This work was supported in part by the National Science Foundation of China (Nos. 62276056 and U24A20334), the Yunnan Fundamental Research Projects (No.202401BC070021), the Yunnan Science and Technology Major Project (No. 202502AD080014), and the Program of Introducing Talents of Discipline to Universities, Plan 111 (No.B16009).
@@ Appendix A Appendix
A.1 Prompt Example
ASR 1. Hear the audio clip and transform it into text format. <|AUDIO|>
2. Analyze the recording and summarize its contents. <|AUDIO|>
Table 4: Examples of prompts for different tasks.
A.2 Details of Models and Datasets
In this paper, we leverage multiple audio encoders and LLM to construct the end-to-end speech LLM. Our training dataset is sourced from commonly used open-source datasets, totalling approximately 450 hours of audio data, corresponding to 313,208 samples, as outlined in Table 5. For SNV, we randomly concatenate individual utterances to form new speech signals with the number of speakers ranging from one to four.
@@ A.2 Details of Models and Datasets
Kim et al. (2019)
Table 5: The whole training dataset.
In our paper, we adopt multiple pre-trained audio encoders and LLMs, and we list the architecture settings for all models we used in our experiments in Table 6. Notably, for the Whisper model, we only used its encoder part as an audio feature extractor.
@@
Table 6: The settings of the pre-trained model we used in our experiments. For the audio encoder models, we utilize only the encoder component and freeze all parameters. For the LLMs, we freeze the base model parameters and apply LoRA adapters to fine-tune the model.
A.3 Details of Baseline Implement
In our work, we compare our method against two types of baselines. The first baseline consists of models using a single encoder, while the second baseline involves fusing multiple audio encoders, either as in previous work Hu et al. (2024); Tang et al. (2024) or through an averaging operation.
For the single encoder baseline, we train the model using the same settings as in our method. For the second baseline, we train the model using the open-source codebases from WavLLM and SALMONN. We integrate the adapter components from these repositories into our code and train the baseline model using our training data, employing the same pre-trained audio encoders and LLMs as in our method. During training, we applied the same hyperparameters as our method. Since we use different encoders and LLMs compared to the baselines, we adjust the dimensions of the adapter to match the specific audio encoder and LLM we used while maintaining other dimensions independent of the audio encoders and LLM unchanged. Notably, we trained the SALMONN with query length=32 (as training with the original setting query length=1 failed) to ensure comparable performance with the other baseline methods.
A.4 Results on Audio Caption with More Metrics
Due to the multiple evaluation metrics for the AC task, we scored the AC and AC QA tasks using more metrics in Table 7, including CIDEr, SPICE (with coco-caption toolkit), FENSE metrics, and Sentence-BERT [^8] We used the FENSE open-source repository GitHub and scored the entire dataset with eval_system.py. For the SBERT model, we loaded the paraphrase-mpnet-base-v2 model, and for the echecker, we used echecker_clotho_audiocaps_base. However, we encountered a bug when loading the echecker model, which had an unexpected key encoder.embeddings.position_ids. To resolve this, we set strict=False. Additionally, we included results based on similarity using the SBERT model.. We found that the different models exhibited similar performance trends across almost all metrics.
Models
AudioCaps AudioCaps QA
[…]
Single-encoder Baselines
   - Whisper 32.96 0.108 0.596 0.431 0.158 15.04 0.105 0.402 0.205 0.083
   - WavLM 27.14 0.106 0.500 0.290 0.122 12.77 0.104 0.337 0.127 0.049
   - Wav2Vec2 23.81 0.096 0.414 0.205 0.095 11.20 0.092 0.282 0.073 0.038
Multi-encoder Baselines
   - WavLLM 34.93 0.109 0.640 0.569 0.175 16.35 0.108 0.448 0.303 0.109
   - SALMONN 34.86 0.109 0.631 0.542 0.158 15.97 0.108 0.432 0.254 0.093
   - Average 33.22 0.108 0.615 0.471 0.166 15.53 0.108 0.425 0.229 0.093
PaM 35.47 0.111 0.644 0.581 0.183 15.70 0.108 0.428 0.267 0.087
Table 7: More results based on various metrics on the AC task. The sBERT represents Sentence-BERT.
A.5 PaM with More Encoders
We added the BEATs encoder to our framework (which includes four encoders) and found that it significantly improves the performance of our system on the AC task (Table 8). Although fusing the new encoder had some effect on the AMI and SNV tasks, incorporating the BEATs encoder improved the system’s average rank across downstream tasks (Table 9). We plan to conduct additional experiments with the EAT encoder Chen et al. (2024a) in further work.
Prompt-aware Mixture (PaM) LibriSpeech AMI SNV AudioCaps AudioCaps QA
[…]
   - PaM 3.65 7.07 12.79 42.8% 35.47 15.70
   - PaM (BEATs) 3.76 7.22 13.11 49.2% 35.70 16.36
Table 8: Results of PaM with more encoders.
Model Whisper WavLM Wav2Vec2 WavLLM SALMONN Average PaM (audio-based) PaM (prompt-aware) PaM (BEATs)
AVG Rank 7.7 7.5 7.3 4.5 5.0 5.3 3.5 2.5 1.7
Table 9: Average result rank in all downstream tasks of different models.
A.6 Details of Training and Inference Parameters
We train our model for five epochs with a learning rate of 5e-5, 2000 warmup steps, and bf16 precision. We freeze all encoders and the LLM, only training adapters and the fusion modules. For the LLM, we apply LoRA Hu et al. (2022) with a rank of 32 and an alpha of 64, adding LoRA only on the q_proj and k_proj. Each task has the same probability during training. During the inference stage, we select the last checkpoint on the validation set and perform greedy search.
A.7 WavLLM and PaM with Larger LLM
We experiment with WavLLM (concatenation) and PaM under a larger scale LLM based on Qwen2.5-7B and Llama3.1-8B, as shown in Table 10. We found that PaM consistently outperforms concatenation on LibriSpeech. However, in noisier ASR scenarios such as AMI, concatenation performs better. On tasks like SNV, AC, and AC QA, concatenation’s performance is not stable. In the Qwen-based speech large language model, the performance on these three tasks is better than PaM, but in the Llama-based model, the performance on these tasks is significantly worse than PaM.
Models
LibriSpeech AMI SNV AudioCaps AudioCaps QA
[…]
Qwen2.5-7B
   - WavLLM 5.13 10.81 13.66 51.1% 36.27 16.55
   - PaM 3.68 08.36 15.26 43.7% 35.46 15.71
LLaMA3.1-8B
   - WavLLM 6.38 14.08 13.95 03.1% 34.91 14.76
   - PaM 4.85 08.87 15.01 50.8% 35.81 15.82
Table 10: Results based on our PaM adapter and WavLLM with larger LLM.
Notably, we note that the concatenation method in the Llama-based model performs significantly worse on SNV, with only 3.1% accuracy. This is because it becomes difficult to follow the instructions of the SNV task during the inference stage. After further experiments, we found that the concatenation method becomes progressively less effective on SNV. This suggests that the method struggles to achieve a balance between multitasking as training progresses.
A.8 Ablation Experiments of Routing Method
We adopt a prompt-aware routing method to better utilize the information in the prompt based on the LLM, as described in Equations 7 and 8. To further evaluate the impact of different routing strategies, we conducted ablation experiments on various forms of routing methods, as presented in Table 11.
  • Audio-based Routing Method. The routing method employed in most of the MoE models, such as Switch Transfomers Fedus et al. (2022), DeepSeekMoE Dai et al. (2024), and MoWE Zhang et al. (2024), which use the current layer input as the routing module input, and optimize routing module directly based on the loss of outputs. However, ignores information from task labels.
    G $=\text{Top-}k(\text{Softmax}(\mathbf{X}))$ (11)
    We set $k=1$ in the $\text{Top-}k$ function, consistent with the configuration used in our PaM setup. In our model, the MoE layer is positioned after multiple encoders, since we use the fused features from multiple encoders as the routing inputs.
    $\mathbf{X}$ $=\text{FFN}(\text{Concat}(\mathcal{H}^{\text{last}}))$ (12)
In addition, we also performed ablation experiments with our PaM routing method.
  • Without Shared Expert. We maintain the full model configuration but remove the shared expert.
  • Without Task Label. We retain the use of task-related information extracted from the LLM prompt as input to the routing module, without any additional labeling information. Specifically, we remove the auxiliary loss term $\mathcal{L}_{\text{G}}(\text{P}(\text{Task}|\mathbf{H}_{\text{prompt}}),\text{Task})$ in Equation 2.4. In contrast to the audio-based routing method, this variant of the PaM routing method uses the prompt feature $\mathbf{X}_{\text{prompt}}$ as input but without the task label.
[…]
MoE (audio-based) 4.12 8.32 11.21 26.00% 35.66 15.61 3.17
PaM (with one expert) 4.27 10.04 13.77 45.80% 35.76 16.47 2.83
PaM (without shared) 4.31 7.89 13.43 45.20% 35.27 16.35 3.33
PaM (without task label) 3.97 7.97 13.14 43.50% 35.54 15.42 3.17
PaM (ours) 3.65 7.07 12.79 42.80% 35.47 15.70 2.50
Table 11: Ablation results on our routing method and the result based on the audio-based routing method
We found that the PaM routing method outperforms audio-based routing on most ASR and AC QA tasks, especially SNV tasks. This suggests that, in our setting, PaM is superior to the basic MoE routing method for fusing multi-encoder features.
For the PaM without shared experts, we found that it still outperforms the single model baseline and maintains better or comparable performance compared to the multi-encoder baseline on almost all tasks. Compared to PaM without shared experts, PaM with shared experts gains on several tasks but is slightly weaker on AC QA and SNV. This suggests that while shared experts may slightly degrade performance on a few tasks, they can significantly improve the overall effectiveness of the PaM model.
Compared to PaM without task labels and PaM, we found that PaM achieved improvement on most of the tasks, which further illustrates the effectiveness of incorporating task information in the prompt when handling multiple downstream tasks.
A.9 Initialization setup for LLM
We train the LLM model from the open-source base model, following the setup used widely in prior work Hu et al. (2024); Ma et al. (2024b); Tang et al. (2024); Rubenstein et al. (2023a). An interesting alternative is to initialize our LLM module with the LLM module in a pretrained multimodal model like Qwen-audio. While we are not adapting such a setup because Qwen-Audio’s feature alignment was specifically designed for Whisper’s encoder, which can potentially "overfit" to features from Whisper. Our new experiments in Table 12 reveal that the feature spaces of Wav2Vec2 and WavLM are relatively similar, while Whisper’s feature space shows greater divergence. This pattern is also reflected in the weight distribution in Figure 3, where Wav2Vec2 and WavLM appear more closely aligned and significantly different from Whisper. Therefore, using other encoders and restructuring the projector module would still require re-adapting the LLM to comprehend new features.
Cosine similarity emb 1 2 3 4 5 6 7 8 9 10 11 12
Whisper & WavLM -0.91 -0.76 -0.79 -0.83 -0.83 -0.85 -0.89 -0.89 -0.89 -0.84 -0.69 -0.38 -0.85
Whisper & Wav2Vec2 -0.90 -0.73 -0.77 -0.80 -0.81 -0.84 -0.89 -0.89 -0.89 -0.85 -0.75 -0.69 -0.86
WavLM & Wav2Vec2  0.99  0.13  0.26  0.44  0.52  0.60  0.95  0.99  0.99  0.52  0.06 -0.38  0.48
Table 12: The cosine similarity of the hidden states between the layers of different encoders.
A.10 All Detailed Results